CARA Telecon Minutes Date: May 21, 2001 Time: 1:00 PM Attendance: BU: Cornell: CMU: Northwestern: G. Novak R.I.T: S.A.O.: T. Stark, A. Lane, South Pole: C. Martin Chicago: J. Carlstrom, N. Odalen R. Landsberg, R. Loewenstein J. Bausch UCB: W. Holzapfel UCSB: J. Ruhl Approval of previous minutes: APPROVED Agenda Items IP Addresses: R. Loewenstein discussed with the group that Raytheon wants to change IP address at Pole. R. Loewenstein is working with Raytheon on understanding their request for a change and how to best handle any change and its effect on CARA. It is R. Loewenstein's understanding that the change will not occur during winter operations. There has been talk by Raytheon that they will not support Apple Talk. IP address change and Apple Talk will be discussed during the SPUC meeting. It should be noted that Jeff Thompson (consultant to NSF/Raytheon) is the individual that is spearheading these changes. Support for the Apple Talk network for Pole is in the SIPs. *** JC: this issue should be discussed during summer Raytheon planning meeting and also at SPUC -- required support for appletalk or other networking should be explicit in our SIP *** News from the Pole: C. Martin stated that all telescopes are on line and plugging away. According to B. Holzapfel - M. Newcomb claims as of yesterday Acbar is running. M. Newcomb has fixed some software bugs and should be running full time. There was only damage to a disk drive during the overheating problems. DASI - J. Carlstrom stated that everything is running well except the backup computer. DASI team could shut down this computer but it is nice to have it keeping duplicate disks. DASI team will have a telecon with B. Reddall to discuss this problem. AST/RO - T. Stark reported that they are plugging away with the HDO at 464 GHz. Next phase is to continue with galactic center CO 7-6 and CI mapping. C. Martin stated that for a brief period tau was at 1.9 this past week, but most of the time tau has been around 1.2. During the 1st week of May while the FTS was running tau was at 1.0 and spectra was taken. R. Chamberlin has been analyzing the FTS data taken during this time frame, but it is not yet done. *** If weather get down below 0.8 then AST/RO will run the FTS. However, due to the short supply of helium, the FTS should not be run on a routine basis anymore. Let's instead try to at use it to get a "best" spectrum that will hopefully show the 200 um window*** Iridium data transmission - R. Loewenstein has been talking to the engineers to check if software could be installed in iridium phones at Pole that would handle the data transmission. It was determined that we can't install the software at Pole. Instead we will need to install a flash box and/or possibly bring the phones out of Pole to have the software installed. R. Loewenstein is having the engineer contact Raytheon. If the data transmission can be accomplished then we can have email 24/7 at Pole. C. Martin wanted to know if the telephone dropout could be fixed. R. Loewenstein thinks that it is satellite problems -more ground stations will be needed to eliminate the dropout. J. Ruhl reported that the SIP listed 34 liters per day is the correct number for instrument liquid helium usage next year, but this number ignores the 100 & 200 liter transport dewars. J. Ruhl will send out a memo. ***J. Ruhl to write a sentence each on 1) what is needed for instruments, 2) what is needed on station, 3) what is needed in dewars (beyond the existing dewars! - the main point is too indicate clearly that the existing dewars are not sufficient at present *** Giles is trying to understand flow meter. Currently flow meter is giving us 20 liters per day boil off rate. Judging from the differential pressure gauge, it is more like 40 liters per day. The differential pressure gauge has has a history of incorrect readings (shows 2/3 full when full), but actually has been a good predictor of when the helium will run out. DPG (differential pressure gauge) had error at McMurdo when dewar was being filled. Weight showed that dewar was full, but DPG was not showing a full dewar. The DPG is going down linearly and... appears it will hit bottom in August. P. Sullivan expressed concern on the plumbing. He will not insure that plumbing is leak proof and therefore the flow readings might be low. ***C. Kaminski to look at pluming to see if leaks are possible. C. Kaminski to observe only.*** 2 weeks from today helium team should have a good view on when helium will run out. Groups were asked to review observing strategy given the possibility that the helium will run out in August. I.e., AST/RO should concentrate now on the projects with the highest scientific return and that will likely be written up in a timely manner. FTS should not be run except when the weather is exceptionally good and we expect to detect the 200 um atmospheric window. ACBAR should concentrate on CMB power spectrum and ensure calibration is in hand.